### v1

# On Nanocones as a Gravitational Analog System

Abstract
This study delves into the fundamental properties of graphene and boron nitride (BN) nanostructures, exploring their torsional energy characteristics within the framework of Teleparallel Equivalent of General Relativity (TEGR). By constructing nanocones with disclination defects in these materials, we investigate the linear dependence of torsional energy on the disclination angle, as predicted by TEGR. The qualitative validation of TEGR’s energy expression is supported by our simulations, which show a strong correlation between the torsional energy and the disclination angle, consistent with the theoretical predictions. Furthermore, we propose a quantitative analysis by estimating the coupling constant $k$ associated with TEGR through molecular simulations and Density Functional Theory (DFT) calculations. Our results suggest that $k$ reflects the interatomic forces within the materials, providing insights into the nature of spacetime and gravitational interactions on a microscopic scale. These findings not only contribute to our understanding of material physics but also offer implications for the precision and validity of TEGR in describing gravitational phenomena.
June 8, 2024
I Introduction
Analog models in physics have long existed. A mathematical equivalence between the equations governing certain physical phenomena allows for analogy between them. A good example of this occurred in the realm of electrical circuits. In the early 19th century, when instrumentation for measuring electrical and magnetic quantities was being developed, it was very costly to conduct investigations in circuits containing capacitors and inductors. Therefore, mechanical analog models were employed, which facilitated a better understanding of electrical circuits, as well as simpler experimentation. Analog models thus enable the testing of physical concepts without the need to access the physical system of interest. Experimental black hole physics is still in its infancy and could greatly benefit from the use of gravitational analog models barcelo . While experimenting with gravitational waves can shed light on certain aspects of black hole physics, testing other concepts such as Hawking radiation requires the use of analog models, often established in condensed matter. However, there are also acoustic analogs; acoustic black holes, proposed by Unruh Unruh , have been used to understand how event horizons can produce a spectrum of thermal radiation Unruh1 . Hence, gravitational analog models have become important tools for understanding black hole physics. One such example was the construction of an experimental apparatus where a robophysical model was built to simulate the motion of particles around a Schwarzschild singularity li2023robophysical , an experiment that culminated in very satisfactory results compared to the expected values from GR.
General Gelativity (GR) is the most widely accepted theory in the literature describing gravitation. However, it is fraught with inconsistencies. Perhaps the most important of these is the definition of gravitational energy. Indeed, there are several attempts to define such an important concept in physics within the scope of general relativity, but none of them meets the necessary requirements for a consistent definition, namely: i) the energy should be independent of the coordinate system used; ii) it should be the component of a covariant four-vector under global Lorentz transformations and therefore should depend on the choice of reference frame. Alternatively, there are approaches that describe the gravitational field, in which the concept of energy is more natural. In this article, an alternative theory dynamically equivalent to general relativity, called Teleparallel Equivalent to General Relativity (TEGR) maluf , will be used. In the context of TEGR, there exists a consistent energy-momentum four-vector, so this quantity can be used to experimentally determine which approach correctly describes gravitation. In other words, from a dynamic point of view, both GR and TEGR predict the same behaviors. On the other hand, a consistent definition of energy exists only in TEGR (for a complete discussion of gravitational energy, we encourage the reader to consult Ref. annalen ). On one hand, there are currently no experiments designed to deal with gravitational energy. On the other hand, the use of gravitational analogue models may lead to clues suggesting the supremacy of TEGR over GR, if gravitational energy is confirmed in such systems.
Graphene is a material predicted in the mid-20th century and synthesized in the early 21st century grafeno . It consists of a two-dimensional lattice of carbon atoms arranged in a hexagonal shape, giving graphene highly desired properties such as good conductivity and strength. The equation governing the dynamics of graphene is relativistic, as the Dirac equation embodies Lorentz transformations. The Dirac equation also naturally couples to the dynamic variables of TEGR (Teleparallel Equivalent of General Relativity), making graphene an excellent analog gravitational system that could allow investigation into the physical reality of gravitational energy.
When studying the propagation of fermionic quasiparticles in the presence of a $U(1)$ gauge theory, Zubkov and Volovik zubkov2015emergent discovered that, in the presence of elastic deformations, the geometry experienced by the particle was the Weitzenböck geometry. The TEGR is a geometrical gravitational theory within the Weitzenböck geometry. Hence, in this article, we aim to simulate a gravitational field as a non-trivial geometry on a graphene sheet produced by deformations. The connection between geometry and deformations in crystals was established by Katanaev in the Geometric Theory of Defects (GTD) katanaev2005geometric , where elastic deformations are characterized by deviations from the metric tensor in Cartesian coordinates, and plastic deformations arise from the emergence of curvature and torsion. Plastic deformations are caused by topological defects, which can manifest as point defects, line defects, or planar defects. Point defects are often considered in Chern-Simons geometric theory (see, for instance, katanaev2020point ; holz1988geometry ) while planar defects do not seem to hold much significance in analogies between gravitation and defects in crystals. Linear defects can be divided into two categories: disclinations and dislocations. Dislocations are associated with translational deformation, while disclinations are associated with rotational deformations. Disclinations were investigated within the GTD in Ref. katanaev2021disclinations and wedge disclinations in Ref. katanaev2003wedge . Additionally, the combination of two types of dislocations was considered in Ref. katanaev2023combined . Several applications of the GTD can be found in the literature, such as in the propagation of elastic waves katanaev2016rotational and calculating the trajectories of phonons without having to deal with complicated boundary conditions de1998geodesics . In this article, we consider a distinct approach: using the notion that plastic deformation on a graphene sheet can be described through the GTD, we employ the tools of TEGR to obtain a behavior for the torsion energy of the defect.
In pursuit of our goal, we consider a graphene sheet as a background flat geometry and impose a gravitational field by inserting a topological defect into the sheet. We choose a disclination defect to construct a well-known nanostructure for graphene, i.e., a nanocone charlier2001electronic ; adhikari2012stabilities ; ardeshana2017approach . Consequently, we evaluate the gravitational energy of the disclination using TEGR and compare the predicted qualitative behavior of the torsional energy with the energy associated with the torsion potential obtained from computational simulations within Universal Force Field (UFF) . The Universal Force Field (UFF) is a versatile molecular mechanics force field used in computational chemistry. It employs empirical parameters to model interatomic interactions, including bonded and non-bonded terms, torsional potentials, and geometric optimizations. UFF is widely applied in molecular simulations for accurate predictions of molecular structures, energies, and properties. The main aspects of this theory necessary for the comprehension of the simulations will be presented in this article. We also infer about the possibility of the physical origin of the coupling constant in graphene and obtain a numerical value for the energy within the order of magnitude given by the simulations. Therefore, we aim to address these aspects in this article to support future experiments designed to measure gravitational energy, especially since the synthesis of nanocones is already a reality experimental .
Thus, this article is divided as follows: In Section II, we briefly present the concepts of gravitational energy and momentum within the scope of TEGR. In section III we introduce the disclination topological defect considered. In Section IV, we present the general aspects of the UFF. In Section V, we present the results and respective analyses. Finally, in the last section, we present our conclusions.
II Teleparallelism Equivalent to General Relativity (TEGR)
Teleparallelism Equivalent to General Relativity (TEGR) is a geometrical theory of gravity that was first considered by Einstein in an attempt to unify gravity and electromagnetism. Several decades after being abandoned by Einstein, Møller revived the idea with another objective: to establish a true energy-momentum for the gravitational field moller1964conservation . Despite the many decades of GR, researchers have not been able to define a true energy-momentum tensor for the gravitational field, with several pseudo-tensors defined in the literature, all of which depend on the choice of coordinates. Therefore, a definition for the energy of a gravitational field is only possible globally, i.e., by integrating the pseudo-tensors across the entire spacetime. Such integration is possible only if the spacetime is asymptotically flat at spatial infinity. After a few years of attempting to define a true energy-momentum tensor for the gravitational field, Møller realized that the main problem with GR was the fact that its fundamental variable is the metric tensor $g_{\mu\nu}$ and concluded that he could solve the problem by formulating GR with the tetrad field $e_{a\mu}$ as its fundamental variable. Furthermore, Maluf realized that the field equations of the theory allow for a definition of a true energy-momentum tensor for the gravitational field. In the following, we follow the development of TEGR as presented in Ref. maluf .
The tetrads $e_{a\mu}$ have 16 independent components, with one $\mu$ (Greek) raised and lowered by the metric tensor of spacetime $g_{\mu\nu}$ and $a$ (Latin) raised and lowered by the metric tensor $\eta_{ab}$ of the tangent space (fiber bundle) to the spacetime at the event $x^{\mu}$. Thus, the tetrads project quantities $v^{\mu}$ from spacetime into $v^{a}$ quantities of the tangent space, i.e., $v^{a}=e^{a}\,_{\mu}v^{\mu}$. Given the properties of lowering and raising indices by the metric tensor, the metric tensor itself can be obtained by the relation
$g_{\mu\nu}=e_{a\mu}e^{a}\,_{\nu}\,.$ (1)
In view of equation (1), the tetrads are more fundamental than the metric tensor. The tetrads $e_{a\mu}$ transform as covariant 4-vectors under coordinate transformations in the Greek index and under the $SO(3,1)$ group in the Latin index.
The geometry in which TEGR is established in the Weitzenböck geometry. Given the prominence in the literature of topological defects considered within Riemann-Cartan geometry, we depart from Riemann-Cartan to the Weitzenböck one, where the Levi-Civita connection is responsible for the rotation symmetry instead of the affine connection in Riemann-Cartan geometry. Therefore, we shall show that TEGR admits linear rotations (disclinations) and translations (dislocations) as topological defects in its geometry, similar to Riemann-Cartan.
In a manifold endowed with metricity, the structure equations can be obtained from the co-frame 1-form $e^{a}=e^{a}\,_{\mu}dx^{\mu}$ and the connection 1-form $\omega^{a}\,_{b}=\omega^{a}\,_{\mu b}dx^{\mu}$, where $\omega^{\lambda}\,_{\mu\nu}=e_{a}\,^{\lambda}e^{b}\,_{\nu}\omega^{a}\,_{\mu b}$ are the 0-form components of the connection. The torsion 2-form $T^{a}$ of the manifold can be obtained by the exterior total derivative $D$ of the co-vector, i.e.,
$T^{a}=D\,e^{a}$ $=d\,e^{a}+\omega^{a}\,_{b}\wedge e^{b}$
$=\frac{1}{2}e^{a}\,_{\lambda}T^{\lambda}\,_{\mu\nu}dx^{\mu}\wedge dx^{\nu}$ (2)
where $d$ indicates the exterior derivative, $\wedge$ denotes the wedge product, and
$T^{\lambda}\,_{\mu\nu}=e^{a}\,_{\lambda}(\partial_{\mu}e^{a}\,_{\nu}-\partial_{\nu}e^{a}\,_{\mu}+\omega^{\lambda}\,_{\mu\nu}-\omega^{\lambda}\,_{\nu\mu})$ (3)
are the 0-form components of the torsion tensor. The curvature 2-form is
$R^{a}\,_{b}$ $=d\omega^{a}\,_{b}+\omega^{a}\,_{c}\wedge\omega^{c}\,_{b}$
$=\frac{1}{2}e^{a}\,_{\alpha}e_{b}\,^{\beta}R^{\alpha}\,_{\beta\mu\nu}dx^{\mu}\wedge dx^{\nu}\,,$ (4)
where
$R^{\alpha}\,_{\beta\mu\nu}$ $=\partial_{\mu}\omega^{\alpha}\,_{\nu\beta}-\partial_{\nu}\omega^{\alpha}\,_{\mu\beta}$
$+\omega^{\alpha}\,_{\mu\rho}\omega^{\rho}\,_{\nu\beta}-\omega^{\alpha}\,_{\nu\rho}\omega^{\rho}\,_{\mu\beta}$ (5)
are the components 0-form of the curvature tensor.
The geometry of TEGR may be understood as a special case of Cartan geometry, i.e., when the curvature tensor (4) constructed from the affine connection vanishes. In this case, the affine connection $\omega^{a}\,_{c}$ plays no role in the dynamics of the field maluf . Hence, we may consider it to be zero. Given the relation of the curvature tensor with rotational symmetries, it is possible to construct a geometry in which the parallel transport of a vector along a closed curve occurs, i.e., a parallelism at a distance ($\tau\eta\lambda\epsilon$ =“tele” = far). Thus, in such a geometry, the tetrad has zero covariant derivative, i.e.,
$\nabla_{\mu}e_{a\nu}=\partial_{\mu}e_{a\nu}-\Gamma^{\lambda}\,_{\mu\nu}e_{\lambda a}=0\,.$ (6)
The connection that guarantees distant parallelism is given by
$\Gamma_{\mu\lambda\nu}=e^{a}\,_{\mu}\partial_{\lambda}e_{a\nu}$ (7)
whose torsion tensor components are given by
$T^{a}\,_{\lambda\nu}=\partial_{\lambda}e^{a}\,_{\nu}-\partial_{\nu}e^{a}\,_{\lambda}\,.$ (8)
From the torsion tensor components, the contortion tensor
$K_{\mu\lambda\nu}$ $=$ $\frac{1}{2}(T_{\lambda\mu\nu}+T_{\nu\lambda\mu}+T_{\mu\lambda\nu})\,,$ (9)
can be evaluated. Since the affine connection is chosen to be zero, the components of the Levi-Civita connection $\mathring{\omega}_{\mu ab}$ can be evaluated using the relation
$\Gamma_{\mu\lambda\nu}=\mathring{\omega}_{\mu\lambda\nu}+K_{\mu\lambda\nu}=0$ (10)
for the contortion tensor components.
After establishing the geometry on which TEGR is constructed, we can define the theory on this geometry. First, we define the superpotential as
$\Sigma^{abc}=\frac{1}{4}(T^{abc}+T^{bac}-T^{cab})+\frac{1}{2}(\eta^{ac}T^{b}-\eta^{ab}T^{c})\,,$ (11)
where $T^{\mu}=T^{\nu}\,_{\nu}\,{}^{\mu}$. Contracting the superpotential with the torsion, we may construct the Weitzenböck invariant $T=\Sigma^{abc}T_{abc}$ and the Lagrangian density of TEGR
$\mathfrak{L}(e_{a\mu})$ $=$ $-\kappa\,e\Sigma^{abc}T_{abc}-\mathfrak{L}_{M}$ (12)
$=$ $-\kappa\,e\,(\frac{1}{4}T^{abc}T_{abc}+\frac{1}{2}T^{abc}T_{bac}-T^{a}T_{a})$
-- $\mathfrak{L}_{M}\,,$
where $e=det(e^{a}\,_{\mu})$, $k=1/16\pi$ is the coupling constant ($k=c^{3}/16\pi G$ in SI units), and $\mathfrak{L}_{M}$ stands for the matter-radiation fields Lagrangian density. The torsion invariant, and hence the Lagrangian density, is equal to the curvature scalar constructed from the Levi-Civita connection plus a total divergence term, i.e.,
$eR(e)$ $=-e(\frac{1}{4}T^{abc}T_{abc}+\frac{1}{2}T^{abc}T_{bac}-T^{a}T_{a})$
$+2\partial_{\mu}(eT^{\mu})\,.$ (13)
The Lagrangian density (12) is invariant under coordinate transformations and under global Lorentz transformations, but not under local Lorentz transformations. The total divergence term $2\partial_{\mu}(eT^{\mu})$ makes (13) invariant under local Lorentz transformations, and its inclusion in the Lagrangian density would make it invariant under such transformations. In this case, we would have another theory than TEGR. However, the equations of motion would not be altered.
The dynamics of the theory can be obtained by varying the Lagrangian density (12) with respect to $e_{a\mu}$, yielding
$\partial_{\nu}\left(e\Sigma^{a\lambda\nu}\right)=\frac{1}{4\kappa}e\,e^{a}\,_{\mu}(t^{\lambda\mu}+T^{\lambda\mu})\,,$ (14)
where $T^{\lambda\mu}$ is energy-momentum tensor of the matter-radiation field and
$t^{\lambda\mu}=\kappa\left[4\,\Sigma^{bc\lambda}T_{bc}\,^{\mu}-g^{\lambda\mu}\,\Sigma^{abc}T_{abc}\right]\,.$ (15)
Although the Lagrangian density (12) is not invariant under local Lorentz transformations, the field equations (14) are covariant under coordinate transformations and under the full group of Lorentz transformations. It can be shown that (14) may be written as $R_{a\mu}-\frac{1}{2}e_{a\mu}R=\frac{1}{2k}T_{a\mu}$, i.e., they are equivalent to Einstein’s equations. Therefore, TEGR yields the same dynamics as GR but allows for the definition of a true energy-momentum tensor.
The energy-momentum tensor can be obtained directly from (14). By noticing that the superpotential is antisymmetric in the last two indices, we have
$\partial_{\lambda}\partial_{\nu}\left(e\Sigma^{a\lambda\nu}\right)=0\,.$ (16)
Integrating (16) over an arbitrary three-volume $V$ and separating the time and spatial coordinates, we arrive at a continuity (or balance) equation
[…]
where $dS_{j}$ is the infinitesimal area element of $\partial V$ oriented along the $j$ direction. If the RHS of (17) is taken at spatial infinity, we have a conservation equation for the quantity
$P^{a}=\int_{V}d^{3}x\,e\,e^{a}\,_{\mu}(t^{0\mu}+T^{0\mu})\,,$ (18)
and we may identify (15) as the energy-momentum tensor of the gravitational field. The tensor (15) is a true tensor, i.e., it transforms covariantly under coordinate transformations. Given this identification, the quantity
$P^{a}=4\kappa\,\int_{V}d^{3}x\,\partial_{\nu}\left(e\,\Sigma^{a0\nu}\right)$ (19)
is the total momentum 4-vector of the gravitational field contained within $V$. The 4-momentum (19) is invariant under coordinate transformations of the three-dimensional space, under time reparametrizations, and covariant under global Lorentz transformations.
III Topological Defects in spacetime
Linear topological defects in spacetime can be constructed using the Volterra process, which distorts spacetime. This process involves cutting and soldering forms that allow a section of spacetime to be removed and the edges glued together. The details of this process can be found in Ref. puntigam1997volterra . Ten defects can be constructed using this method: four temporal ones, three spatial dislocations, and three spatial disclinations.
A dislocation is characterized by the Burgers vector $b^{(i)}$ oriented along the $(i)$ directions, representing the translational deficit in the topology. The Burgers vector can be evaluated from the torsion tensor using the relation katanaev2005geometric
[…]
where $S$ is the surface orthogonal to the defect. Hence, from (20), we can see that the torsion tensor represents the density of dislocations in space. If the integral (20) along a surface $S$ oriented by $dx^{\mu}\wedge dx^{\nu}$ evaluates to a nonzero value, we can conclude that space has a dislocation defect. Applying Stokes’ theorem, we can write (20) as
[…]
i.e., the defect is “measured” by an observer carried along the boundary of $S$ as the closing failure of the closed Burgers circuit $\partial S$. This result is expected, as torsion is associated with translational symmetry. Therefore, given that the Weitzenböck geometry allows a nonzero torsion tensor, it is expected that TEGR allows for the characterization of dislocation defects.
A disclination defect is associated with angular deficits in spacetime and can be evaluated from the matrix katanaev2005geometric
[…]
The orientation of the angle deficit axis can be obtained by the path-ordered exponential of $\Omega^{(i)(j)}$, i.e.,
$\mathcal{G}^{(i)(j)}=\exp{(2\pi\,\Omega^{(i)(j)})}\,.$ (23)
The matrix $\mathcal{G}^{(i)(j)}$, called Frank matrix, represents the rotation matrix along the axis containing the angle deficit, i.e., the rotation angle represents the amount that we must rotate to perform a full $2\pi$ lap in spacetime. The curvature tensor of the affine connection is zero for the Weitzenböck geometry. However, the curvature tensor of the Levi-Civita connection $\mathring{\omega}_{\mu ab}$ is nonzero in the general case. This curvature tensor allows for a consistent definition of disclination in TEGR. This feature will be demonstrated here for the defect considered.
A conical defect along the $z$ direction is characterized by the deficit angle parameter $\Delta\phi=2\pi\big{(}1-\beta(r)\big{)}$ in the line element
$ds^{2}=-dt^{2}+dr^{2}+\beta(r)^{2}r^{2}d\phi^{2}+dz^{2}\,.$ (24)
When $\beta=1$, we recover Minkowski spacetime. The deficit angle can be evaluated from (24) by computing the length of the circumference along the $z$ axis, but we will consider relations (22) and (23). First, we need a set of tetrads associated with the metric (24) and adapted to a static observer. One set satisfying the condition is given by
$e_{a\mu}=\left(\begin{array}[]{cccc}-1&0&0&0\\ 0&\cos\phi&-r\sin\phi\,\beta(r)&0\\ 0&\sin\phi&r\cos\phi\,\beta(r)&0\\ 0&0&0&1\\ \end{array}\right)\,.$ (25)
From tetrads (25), we evaluate the only non-zero component of $T^{\lambda\mu\nu}$ as
$T^{221}=\frac{1-\partial_{r}(r\,\beta)}{r^{3}\beta^{3}}\,,$ (26)
and also the only non-zero component of the Levi-Civita connection as
$\mathring{\omega}_{2}\,^{(1)(2)}=1-\partial_{r}(r\,\beta)\,.$ (27)
From (27), the only non-zero component of $R^{ab}\,_{\mu\nu}$ is given by
$R^{(1)(2)}\,_{21}=2\,\partial_{r}\beta+r\,\partial^{2}_{r}\beta\,.$ (28)
We can see that the curvature tensor is non-zero. The Frank matrix can be evaluated from (22) directly, or we may notice that if the rotations are contained in a spatial plane orthogonal to the defect, as is the case considered here, we have the subgroup $SO(2)$ where the quadratic components of the connection vanish, and we may write (22) as
[…]
Hence,
[…]
The evaluation of (III) involves modeling the defect core, i.e., choosing a function $\beta(r)$. A simple model can be constructed by considering two regions: (i) a disclination core, where we have a flat spacetime, and (ii) an outside region, where we also have a flat spacetime, i.e., both regions are locally flat. In this case, we may express the function $\beta(r)$ as
$\beta(r)=\begin{cases}\beta_{0}\,,&r>a\,,\\ 1\,,&r\leq a\,,\end{cases}$ (31)
where $a$ is the dislocation core radius. We can see that the curvature tensor will be zero everywhere but at $r=a$. Therefore, the spacetime is locally flat but not globally. By choosing a circuit $\partial S$ with radius $R>a$, we have
$\mathcal{G}^{(i)(j)}$ $=exp\left[\left(\begin{array}[]{cc}0&2\pi(1-\beta_{0})\\ -2\pi(1-\beta_{0})&0\\ \end{array}\right)\right]$ (34)
$=\left(\begin{array}[]{cc}\cos(2\pi(1-\beta_{0}))&\sin(2\pi(1-\beta_{0}))\\ -\sin(2\pi(1-\beta_{0}))&\cos(2\pi(1-\beta_{0}))\\ \end{array}\right)\,.$ (37)
i.e., the rotation matrix of the deficit angle $\Delta\phi=2\pi(1-\beta_{0})$ along the $z$ axis. Thus, TEGR allows torsion and curvature and is consistent in computing topological defects in spacetime, with the rotational defects given by the Levi-Civita connection constructed out of the contortion tensor given by (10).
The 4-momentum (19) of (25) has already been obtained in the literature maluf1997gravitational ; maluf2001space ; carneiro2020quantization . However, from the torsion (22), it is straightforward to obtain the only non-zero component of $\Sigma^{a0\nu}$ as
$\Sigma^{(0)01}=\frac{1-\partial_{r}(r\beta)}{2r\beta}\,.$ (38)
Hence, the energy contained within a cylinder of radius $R>a$ and length $L$ is given by
$P^{(0)}$ $=2k\int_{V}\partial_{r}\big{(}1-\partial_{r}(r\beta)\big{)}dV$
[…]
$=4\pi\,k\,L(1-\beta_{0})$
$=2k\,L\,\Delta\phi\,.$ (39)
From (39), we see that with TEGR, we expect the energy to not depend on the radius $R$ and, more importantly, to be linear with the angle deficit $\Delta{\phi}$.
Our goal is to construct a graphene monolayer that has the topological structure of (24), calculate the energy of the structure associated with its molecular torsion potential, and determine the qualitative validity of (39) and TEGR itself.
IV Universal Force Field in Molecular Mechanics
The complexity of multi-electron atoms makes it challenging to apply quantum mechanics directly to solve their problems. In such scenarios, we must resort to approximate methods alongside the many-electron wave function trial $\psi\left(\vec{r_{1}},\vec{r_{2}},\vec{r_{3}},....,\vec{r_{n}}\right)$ bretonnet2017basics ; viana2004quantitative to solve the appropriated Schrodinger equation, and therein obtain the behavior of these systems. Due the computation limitations, molecular mechanics, a classical approach, becomes indispensable for large atomic systems, aiding in understanding molecular properties and behavior in diverse settings. At the core of molecular mechanics lies the Universal Force Field (UFF), a framework crucial for simulating molecular interactions rappe1992uff ; jaillet2017uff .
The UFF is established empirically, parameterizing potential energy functions based on data mackerell2004empirical ; bradley1972ap . It simplifies by focusing solely on nuclear positions, neglecting electron movements, and computes an effective potential through the sum of atomic contributions potentials tajkhorshid2002control ; leach2001molecular , i.e., treats the atoms as particles. The force field encompasses various atomic interactions, including bond stretching, bending, rotations, cross-interactions, and non-bonded interactions. The bonded potential $V_{bonded}$ comprises terms for bonded pairs, valence angles, bending, consecutive stretching, consecutive bending, and cross-interactions. Conversely, $V_{non\,bonded}$ captures non-bonded interactions between atom pairs.
The UFF expression (42) combines these terms into a comprehensive potential energy function, while the torsional potential $V_{torsion}$ (43) handles torsion terms of the combined atoms. These formulations enable detailed simulations of molecular dynamics, essential for exploring the behavior of complex systems. The Universal Force Field is comprised of two general terms given by
$V_{\vec{r_{N}}}=V_{bonded}+V_{no\,bonded}$ (40)
for a system with N nuclei. The potential $V_{bonded}$ is given by
$V_{bonded}$ $=\sum_{d}V(d)+\sum_{\theta}V(\theta)+\sum_{\phi}V(\phi)+\sum_{\chi}V(\chi)$
$+\sum_{b,b^{{}^{\prime}}}V(b,b^{{}^{\prime}})+\sum_{\theta,\theta^{{}^{\prime}}}V(\theta,\theta^{{}^{\prime}})+\sum_{b,\theta^{{}^{\prime}}}V(b,\theta^{{}^{\prime}})$
$+\sum_{\theta,\theta^{{}^{\prime}}}V^{{}^{\prime}}(\theta,\theta^{{}^{\prime}})+\sum_{{\chi},\chi^{{}^{\prime}}}V(\chi,\chi^{{}^{\prime}})\,.$ (41)
The potential $V_{no\,bonded}$ is
$\sum_{i=1}^{N}\sum_{j=i+1}^{N}[V_{Coulomb}+V_{Lennard-Jones}]\,.$
Finally, the Universal Force Field is described by the potential
$V_{\vec{r_{N}}}$ $=\sum_{b}D_{b}\big{[}1-\exp^{-a(b-b_{0})}\big{]}^{2}+\sum_{\theta}H_{\theta}(\theta-\theta_{0})^{2}$
$+\sum_{\theta}H_{\theta}(\theta-\theta_{0})^{2}+\sum_{b,b^{{}^{\prime}}}F_{bb^{{}^{\prime}}}(b-b_{0})(b^{{}^{\prime}}-b_{0}^{{}^{\prime}})$
$+\sum_{\theta,\theta^{{}^{\prime}}}F_{\theta\theta^{{}^{\prime}}}(\theta-\theta_{0})(\theta^{{}^{\prime}}-\theta_{0}^{{}^{\prime}})+\sum_{b,\theta}F_{b\theta}(b-b_{0})(\theta-\theta_{0})$
$+\sum_{i=1}\sum_{j=i+1}\Bigg{(}4\epsilon_{ij}\Bigg{[}\Bigg{(}\frac{\sigma_{ij}}{r_{ij}}\Bigg{)}^{12}-\Bigg{(}\frac{\sigma_{ij}}{r_{ij}}\Bigg{)}^{6}\Bigg{]}+\frac{q_{i}q_{j}}{2\pi\epsilon_{0}r_{ij}}\Bigg{)}$
$+V_{torsion}\,.$ (42)
The first sum in the UFF potential function pertains to bonded atomic pairs, accounting for the interactions between atoms directly linked by chemical bonds. The second term encompasses all valence angles within the molecule, defined by the angle between three atoms A-B-C where A and C are bonded to B. This angle follows a harmonic potential model, reflecting the natural tendency of bonded atoms to maintain specific angles. Moving on, the third term addresses bending interactions, capturing the propensity of atoms to deviate from their ideal bond angles. The fourth term sums up consecutive stretching terms, representing the energy associated with stretching bonds between pairs of atoms. Similarly, the fifth term focuses on consecutive bending interactions, highlighting the energy changes when atoms deviate from their ideal bond angles successively. The third to sixth terms encompass cross-interaction phenomena, such as cross-terminus, out-of-plane tension, quadratic stretching, quadratic bond-angle interaction, quadratic angle-angle interaction, and out-of-plane tension interaction. These terms capture complex molecular behaviors arising from combinations of stretching, bending, and torsional motions, as well as interactions involving multiple bonds and angles. Finally, the last term accounts for non-bonded interactions, considering all pairs of atoms (i and j) in the molecule. This term encapsulates electrostatic interactions (Coulombic) and van der Waals forces (Lennard-Jones), among others, which contribute to the overall energy of the molecular system leach2001molecular .
The torsional potential, denoted as $V_{torsion}$, is expressed as
$V_{torsion}$ $=\sum_{\phi}H_{\phi}\big{[}1+s\,cos(n\phi)\big{]}$
$+\sum_{\chi}H_{\chi}\chi^{2}+\sum_{\theta,\theta^{{}^{\prime}}}F_{\theta\theta^{{}^{\prime}}\phi}(\theta-\theta_{0})(\theta^{{}^{\prime}}-\theta_{0}^{{}^{\prime}}cos(\phi))$
$+\sum_{\chi,\chi^{{}^{\prime}}}F_{\chi\chi^{{}^{\prime}}}\chi\chi^{{}^{\prime}}\,.$ (43)
In equation (43) first term represents the sum of all torsion terms. The second term represents the proper and improper torsion terms (symmetrical and anti-symmetrical). The third term represents 2 bending and a torsion, and the fourth term represents the sum of the terms that include 1 bending and a torsion.
The development of UFF was a significant leap forward in creating a versatile and comprehensive force field because it uses a combination of experimental data and high-level quantum mechanical calculations to parameterise the force field for different types of atoms and their interactions mayne2013rapid ; vanommeslaeghe2014molecular . One of UFF’s major features is the use of atomic hybridization states to refine the parameterization process, which allows the force field to adapt to different bonding environments within the same element.
The versatility of UFF makes it suitable for a wide range of applications in molecular modeling. This has significant implications for fields such as materials science, biochemistry and pharmacology mackerell2004empirical . For example, in drug design, UFF can be used to model the interaction between a drug molecule and its target protein, providing information on the binding affinities and conformational changes involved tute1995drug . In materials science, UFF facilitates the study of molecular structures and properties of polymers, nanomaterials and complex alloys. By accurately modeling the forces within these materials, researchers can predict mechanical properties, stability and reactivity, guiding the design of new materials with customized properties.
The wide applicability and parameterization of the Universal Force Field for almost all elements are among its greatest strengths. This universality makes it a valuable tool for researchers working with diverse chemical systems without the need to switch between different force fields. However, this generality also introduces certain limitations. As the UFF is widely applicable, it may not have the precision of specialized force fields tailored to specific types of molecules or interactions.
In our molecular mechanics calculations for this study, we utilized the Universal Force Field (UFF) within the computational framework provided by the Forcite code rappe1992uff . This approach employs classical mechanics, adhering to Newtonian principles with an empirical potential function to describe molecular interactions. The Forcite code executes calculations by iteratively minimizing the potential energy, a process crucial for achieving a stable system. The iterative procedure aims to reach the global minimum of the potential energy surface, indicating a state where the forces on the system balance out to zero. This self-consistency cycle persists until the convergence criteria are met, typically after multiple iterations. In order to ensure accuracy, we utilized an integration grid to finely sample the potential energy landscape during geometric optimizations. Geometric optimization involves adjusting the atomic coordinates iteratively to locate the configuration that corresponds to the minimum energy state. Our convergence criteria included limiting the maximum atomic displacement to $5.0\times 10^{-3}$ angstroms, the maximum energy change to $1\times 10^{-4}$ kcal/mol, the maximum force to $1.0\times 10^{-3}$ Hartree per angstrom, and the maximum stress to 0.015 gigapascals. These stringent criteria were essential for achieving precise and reliable results in our calculations, ensuring that the molecular configurations obtained were energetically favorable and physically meaningful.
V Results and analysis
The first step in corroborating result (39) is to choose a material that easily emulates the three-dimensional space section of spacetime. The Volterra process can be performed on a nanoscale in a graphene sheet. Graphene is a very stable material with strong carbon-carbon bonds (Young’s modulus of $1TPa$) that can accommodate disclination defects with minimal local deformations for specific disclination angles. This feature is important for properly describing the material by the line element (24), where, for function (31), there must be no local deformations, i.e., the resulting geometry must be locally flat.
The total energy of the structure has several contributions, such as those associated with the terms in equation (43). Part of this energy, i.e., the one associated with the torsion potential, has geometric origins and it is global, meaning it is linked to the overall shape of the structure. This energy is caused by the disclination defect and is not expected to be significantly affected by other atomic phenomena, such as the number of atoms in the structure. Given that the potential from the Universal Force Field (UFF) is empirical, we can regard its results as quasi-experimental. Our goal is to identify this “energy derived of experimental data” with the torsion energy described by TEGR.
In this section, we consider disclination defects in graphene sheets in subsection V.1 and in boron nitride sheets, graphene-like sheets with boron and nitrogen atoms alternating positions, in subsection V.2. Energy will be measured in electron volts (eV) and distances in angstroms ($\AA$ = $10^{-10}$ meters) unless otherwise stated.
V.1 Graphene nanocones
A graphene sheet is not a strictly two-dimensional material, as it has a small but non-zero thickness. This thickness allows the description of the metric tensor of a non-deformed sheet using the metric tensor of the three-dimensional $t=\text{constant}$ hypersurface of Minkowski spacetime.
The sheets are composed of carbon hexagons with partial single-double bonds (aromatic rings). This hexagonal structure of the sheet allows it to accommodate global disclinations only with specific discrete values given by integer multiples of $60^{\circ}$, i.e.,
$\Delta\phi=n\frac{\pi}{3}\,,$ (44)
where $n=1,2,3,4,5$. A graphene sheet with a disclination topological defect characterized by a disclination angle given by (44) is called a nanocone. It has been found that nanocones of silicon carbide (with a similar structure) are stable only for $n=1,2,5$ alfieri2011structural . However, by constructing pentagons at the top of the silicon carbide nanocone, dangling bonds are avoided, and optimized structures can also be constructed for $n=3,4$ adhikari2012stabilities . Therefore, to accommodate as much data as possible, we shall consider structures with a flat top containing one or more carbon pentagons, as silicon carbide is also hexagonal. This approach may make the disclination core less precise in describing the actual defect, but it is necessary to achieve stable structures and avoid local bending. Consequently, we expect a minor error source due to the deviation of the disclination core from a perfect circle described by (31).
The second step is to construct the Volterra process for the graphene sheet. We consider distinct sheets of various radii that are nearly circular. Given the hexagonal symmetry, it is impossible to achieve a perfectly circular sheet, so we anticipate another minor error due to the deviation from a perfect circle of radius $R$ in the integration contained in (39). To construct the nanocone, we cut a sheet, as demonstrated for Sheet 1 (with radius $2R_{1}=48.765\AA$), for an angle of $60^{\circ}$, as shown in Figure 1.
Figure 1: Graphiccal representation of the removal of a section $\Delta\phi=60^{\circ}$ from graphene Sheet 1.
Next, we solder the edges of the sheet. The resulting nanocone is presented in Figure 2.
Figure 2: Nanocone obtained from graphene Sheet 1 with disclination angle $\Delta\phi=60^{\circ}$.
Before simulating, we adjust all the border atoms to ensure that their valence is complete and no atom is left bonded to only one other atom, i.e., all carbon atoms are in aromatic rings before soldering.
[…]
Figure 3: Energy associated with the torsion potential (in eV) for the nanocones as a function of the disclination angle deficit $\Delta\phi$. The dots indicate the data from the simulation, and the lines represent the linear fit.
We can observe an almost linear behavior. By performing a linear fit $E=\alpha\Delta\phi+E_{0}$ of the data, we obtain for Sheet 1
$E_{1}=6.73327\,\Delta{\phi}+0.924306$ (45)
with a correlation of $0.994$.
In order to strengthen our results, we considered three additional sheets, namely Sheet 2 (with radius $54.5602R_{2}=$ 54.560 Å), Sheet 3 (with radius $68.8862R_{3}=$ 68.886 Å), and Sheet 4 (with radius $98.3802R_{4}=$ 98.380 Å), and performed the same procedure, i.e., four nanocones for each sheet. Thus, a total of 16 simulations were conducted for graphene sheets. The results can also be seen in Figure 3. The linear fits are
$E_{2}$ $=7.377\,\Delta{\phi}+0.501\,,$ (46)
$E_{3}$ $=8.854\,\Delta{\phi}-0.574\,,$ (47)
$E_{4}$ $=10.516\,\Delta{\phi}-1.858\,,$ (48)
for Sheets 2, 3, and 4, respectively. The correlations for the linear fits are $0.996$, $0.998$, and $0.991$, respectively.
TEGR predicts an energy independent of the radius, characterized by a universal constant (specific to the material) multiplied by the disclination angle. This universal constant likely depends on an unknown coupling constant $k$ and the thickness of the sheet $L$. Since we do not know the exact value of constant $k$, we can only rely on qualitative corroboration for equation (39). Further discussion on the nature of this constant in graphene and BN will be presented in Section VI.
The findings illustrated in Figure 3 reveal a clear linear relationship between energy associated with the torsion potential and the disclination angle, aligning closely with the predictions of equation (39). This alignment leads us to view these results as providing qualitative support for our theoretical framework, and as a qualitative validation of TEGR.
Upon analyzing equations (46)–(48), we notice a non-zero energy associated with the torsion potential when the flat sheet should ideally have zero energy according to equation (39). We have verified that all graphene flat sheets exhibit zero energy associated with the torsion potential in the simulation. Therefore, considering that the constant terms in equations (46)–(48) are one order of magnitude smaller than the linear coefficient, we disregard these constants as errors.
The potential causes for the deviation of the linear factor, particularly its dependence on the radius of the original sheet, will be discussed in Section VI.
V.2 Boron nitrite nanocones
The significance of the results observed in the graphene nanocone might initially seem specific to graphene alone. However, we posit that the torsional energy expression derived from TEGR is a universal formulation applicable not only to graphene sheets but also to other materials, and even spacetime. In order to improve this assertion, we conducted a similar analysis on another material.
We generated boron nitride (BN) sheets, known for their structural and thermal properties similar to graphene aldalbahi2015variations . Boron nitride possesses insulating properties and exhibits a white coloration. Unlike graphene, boron nitride sheets maintain strength with increasing layers and boast a high Young’s modulus ($0.865TPa$) falin2017mechanical . These characteristics make boron nitride suitable for forming stable nanocones.
The steps involved in constructing nanocones from BN sheets mirror those for graphene and will not be elaborated on here; readers can refer to the previous subsection for details.
For the BN nanocones, we selected six initial sheets of varying radii: Sheet BN 1 (radius $46.7652R_{1BN}=$ 46.765Å), Sheet BN 2 (radius $54.5602R_{2BN}=$ 54.560Å), Sheet BN 3 (radius $63.3542R_{3BN}=$ 63.354Å), Sheet BN 4 (radius $70.1482R_{4BN}=$ 70.148Å), Sheet BN 5 (radius $83.1382R_{5BN}=$ 83.138Å), and Sheet BN 6 (radius $98.7272R_{6BN}=$ 98.727Å). The outcomes of the 24 simulations are presented in Figure 4.
Figure 4: Energy associated with the torsion potential (in $eV$) for the BN nanocones as a function of the disclination angle $\Delta\phi$. The dots represent simulation data, and the lines depict the linear fit.
Once again, the energy associated with the torsion potential exhibits an almost linear growth with the disclination angle across all six structures. This linear trend is quantified through linear regression on Figure 4, yielding the equations
$E_{1BN}$ $=6.456\,\Delta\phi+1.392\,,$ (49)
$E_{2BN}$ $=7.166\,\Delta\phi+0.858\,,$ (50)
$E_{3BN}$ $=8.017\,\Delta\phi+0.051\,,$ (51)
$E_{4BN}$ $=8.314\,\Delta\phi+0.100\,,$ (52)
$E_{5BN}$ $=9.954\,\Delta\phi+0.461\,,$ (53)
$E_{6BN}$ $=9.226\,\Delta\phi+0.164\,,$ (54)
along with their respective correlations: $0.992$, $0.995$, $0.996$, $0.996$, $0.981$, and $0.950$ for Sheets 1 to 6, respectively. All correlations are close to unity.
In equations (49-54), we also note a small constant term, albeit with a significantly smaller magnitude. Furthermore, we observe a linear increase in the coefficient $\alpha$ with the radius of the structure but, like for graphene, not in a linear manner. Sheet BN 1 had a radius less than half of Sheet BN 6, yet this difference did not affect the ratio of the linear coefficients in the same magnitude.
Ergo, we observe that the linear behavior of the energy, associated with the torsion potential, with disclination angles does not seem to be confined to a single material.
VI Final considerations
In this article, we delve into the analogies between material physics and the structure of spacetime. Several analogies between condensed matter physics and gravitation can be found in the literature kraus2019unexpected ; fagnocchi2008analog . We were able to construct a topological defect in spacetime and an equivalent one on graphene and BN monolayers. From the energy definition of TEGR, we predicted a linear response to the disclination angle in the total energy of spacetime. By identifying this total energy with the geometrical (torsional) energy of the nanocones, we found that the definitions of TEGR can be used to evaluate the torsional energy of the material if the constant, analogous to the gravitational coupling one, of the material is determined. We conjecture that these results are not coincidental but part of a greater result, i.e., the energy expression (39) is valid not only for sheets of any material but for all material structures. This conjecture, together with the known analogies between matter and gravity, is a strong indication that the description of gravity by TEGR is more precise than GR. Since the dynamics of TEGR are equivalent to those of GR, all dynamical analogies between GR and material science are valid in TEGR. However, the absence of a consistent definition of gravitational energy in GR makes the results presented here impossible to obtain in Einstein’s theory. Therefore, the results of this article strongly favor TEGR. Given that graphene nanocones are not merely theoretical structures, since they have already been synthesized experimental , further experimental investigations could verify the present theoretical results.
The main results presented in Figures 3 and 4 show small deviations from a perfectly linear behavior independent of the size of the structure. It is worth noting that the Universal Force Field theory presented in section IV is a good approximation compared to experimental results, i.e., it contains several classical approximations to a quantum problem. Without these approximations the evaluation would be virtually impossible for huge molecular systems like the nanocones. Besides, it is a numerical solution, and small deviations are to be expected. Hence, similar to experimental data, simulation data contains an uncertainty that cannot be removed. In addition to these, we have four other causes for the errors.
First, the model constructed in section III considers a circular disclination core with radius $r=a$. In the limit $a\rightarrow\infty$, we have a disclination core along the $z$ axis only, i.e., a cosmic string. In the real case of the structures, the hexagonal pattern of its atoms makes it impossible to have a punctual disclination core, even a circular one with a small radius. In order to mitigate these effects, we considered structures with an intermediate size, avoiding structures with a small radius. For structures with a diameter around eight hexagons, the correlation to the linear behavior was smaller. The size of the structure does not carry out to the core; thus, with intermediate structures, the contribution of these regions is smaller in comparison. One may think that this contribution is the same for the same sheet, but in reality, the configuration of the disclination core changes with the disclination angle. Let us consider Sheet 1, for example. The core(peak) of the nanocone for $\Delta\phi=60^{\circ}$ can be seen in Figure 5.
Figure 5: Disclination core of the nanocone for Sheet 1 with $\Delta\phi=60^{\circ}$.
Here, we observe a pentagon in the disclination core. When $\Delta\phi=120^{\circ}$, a square appears on top, as illustrated in Figure 6.
Figure 6: Disclination core of the nanocone for Sheet 1 with $\Delta\phi=120^{\circ}$.
For $\Delta\phi=180^{\circ}$ and $\Delta\phi=240^{\circ}$, as shown in Figures 7 and 8 respectively, we observe three and four pentagons forming the core.
Figure 7: Disclination core of the nanocone for Sheet 1 with $\Delta\phi=180^{\circ}$.
Figure 8: Disclination core of the nanocone for Sheet 1 with $\Delta\phi=240^{\circ}$.
The border of the disclination core not only changes with the angle but also its size. These distinct cores are necessary to maintain local flatness and minimize the second source of error.
Secondly, we evaluate the energy of a global defect, where spacetime is locally flat. Therefore, the nanocone’s sides must not bend to avoid additional, untraceable sources of energy. While the choices (44) minimize local bending, it does not eliminate it. For very large structures, bending occurs near the borders, deviating from metric (24). This source of error restricted us from choosing very large structures due to bending and very small structures due to the disclination core deviation from circular symmetry.
Thirdly, in our spacetime model, we considered a continuous and infinite spacetime, which is not the case for the sheets. They have a finite radius, resulting in a border where atomic bonds differ from the rest of the structure. These loose atoms increase with the initial sheet’s size, potentially affecting the structure’s properties unpredictably, and causing the energy to vary with the radius. Also, since the sheets are not perfect circles, the integration contour in (39) cannot be a perfect circle with radius $r=R$ in reality, as can be seen in Figure 2. As the approximation to a circle improves with increasing size, we expect this effect to vary with the initial sheet’s size, influencing the linear coefficient between structures. While we attempted to normalize the energy with the area or perimeter of the initial sheet, we found no satisfactory results. Normalizing with the square root of the radius seemed to approximate the linear coefficients, but we couldn’t explain this and left it unaccounted for.
Finally, the Universal Force Field (UFF) is classical, meaning they is non-relativistic. As a result, the UFF cannot accurately represent materials under relativistic dynamics without introducing errors. This limitation is inherent in its design, as it does not account for relativistic effects.
Given the significance of the qualitative results, we are prompted to ask: how can we quantitatively validate expression (39)? This validation hinges on the nature of the coupling constant $k$. If we interpret spacetime as a material, this constant is associated with some of its internal properties. In SI units, $ck$ has the dimension of force. To associate it with the sheet, we must find an internal characteristic of the material represented by $ck$. The Young modulus of the material could partially express this constant, but it yields a magnitude order distinct from the simulation results. Another possibility is the interatomic force between the atoms in the structure, which already has the right dimension. Finding this force is challenging, but we can evaluate the bound energy between the atoms, i.e., the atom-atom electronic interactions. With the bound energy, we can estimate the force by dividing it by the interatomic mean distance. This yields a rough estimate for the force, albeit with the correct magnitude order.
In order to conduct the simulation, we employed another molecular theory, Density Functional Theory (DFT). DFT is an ab initio theory that allows for the study of molecular dynamics by describing the electrons as an electronic density field. Unlike the Universal Force Field (UFF), it is a quantum approach to the problem. Since the theoretical background of this theory is extensive and these simulations are not the main results of our article, we refer the reader to the literature for detailed information parr1983density ; gross2013density .
Initially, we considered a section with 400 atoms in an infinite graphene sheet. The simulation provided a binding energy of $-3176.27\,eV$. In a graphene sheet, each carbon atom forms three aromatic bonds with adjacent carbon atoms, but each bond is shared between two atoms. Therefore, the total number of bonds can be calculated using the formula
Total number of bonds=3×Number of $\text{Total number of bonds}=\frac{3\times\text{Number of atoms}}{2}\,.$ (55)
For a sheet with 400 carbon atoms, we have
Total number of $\text{Total number of bonds}=\frac{3\times 400}{2}=600\,.$ (56)
Thus, there are 600 covalent aromatic bonds in a graphene sheet with 400 carbon atoms. Dividing the binding energy by 600 yields the energy per bond
$\epsilon_{C}=-5.294\,eV\,.$ (57)
Given the distance $d=1.42\AA$ between each pair of atoms, the estimated force between atoms is
$F_{C}=-\epsilon/d=3.733\,eV/\AA\,.$ (58)
From (39), the predicted linear coefficient $\alpha$ is
$\alpha_{C}=2kL=10.452\,eV\,,$ (59)
where $L=140\,pm$ (approximate diameter of a carbon atom) is the thickness of the sheet. We can compare (59) with the results of the linear fit (45-48). This rough calculation provides a very good approximation with the correct magnitude order and even a close numerical match for Sheet 4, where we obtained $\alpha=10.516\,eV$. Notably, larger sheets yielded more accurate results.
Secondly, for boron nitride (BN), a similar analysis yields an energy of
$\epsilon_{BN}=-4.700\,eV\,.$ (60)
In this case, the distance between atoms is $1.45\,\AA$, and the diameter of boron (the larger atom) is $150\,pm$. Hence, we obtain
$F_{BN}=3.232\,eV/\AA\,,$ (61)
and
$\alpha_{BN}=2kL=9.965\,eV\,.$ (62)
Again, we observe the same magnitude order and a very good numerical approximation for Sheet BN 6, where $\alpha=9.226\,eV$. Additionally, we expect lower general values for the coefficient in the BN sheet, as the force between atoms is slightly smaller, reflected in the smaller values in equations (49-54).
If the preceding hypothesis is accurate, as substantiated by these findings, and we conceptualize spacetime as a cellular structure comprised of minuscule particles, we can deduce that the force between these particles corresponds to the coupling constant. Expressed in SI units, this equation results in $c\,k=\frac{c^{4}}{16\pi G}=2.41596\times 10^{42}N$. The immense magnitude of this force indicates that spacetime possesses an extremely high resistance to deformation (about $10^{30}$ times greater than that of a one-square-meter graphene sheet). Consequently, gravity, which arises from the deformation of spacetime, is a remarkably feeble phenomena if the force of deformation is insufficient, meaning that smaller amounts of energy cannot cause considerable deformation of spacetime. This aligns with our understanding of gravity.
Acknowledgements
The authors are grateful to Professor J. W. Maluf for his invaluable assistance in revising the manuscript and for providing insightful ideas that greatly contributed to this work.